2) Find p(A) where p(x) = x2 - 2x + 6, and A =Please show all work
Answer #1

POL) = 6 - 28+6 , A= 6 DA) = A²-2A + 6I = 11ā Vi 51 f-14 15 1. 12 io) Paol 1-3 -9